OpenStudy (anonymous):

(y-2)^2/4 - (x+3)^2/9 =1 HELP Please.. Tell me what kind of conic section this is..

OpenStudy (anonymous):

\[\frac{ (y-2)^{2} }{ 4 } - \frac{ (x+3)^{2} }{ 9 } = 1\]